# Graph Laplacian-based Bayesian Multi-fidelity Modeling

### v1

@@ 5 Numerical results
In this section we quantify the performance of the multi-fidelity approach for a suite of diverse problems in computational physics. These problems are governed by different physical models that include the equations of linear elasticity, Darcy’s flow, Euler-Bernoulli beam theory, and the incompressible Navier Stokes equations. The data types considered include vectors of quantities of interest (QoIs), and one- and two-dimensional fields, where the dimension of the data space, $D$, ranges from $5$ to $10,201$. The size of the low-fidelity datasets ranges from $3,000$ to $10,000$, and in each case the fraction of the number of high-fidelity data points to the number of low-fidelity data points is small; it lies between $0.5\%$ and $3.3\%$. These attributes are summarized in Table 3. Each of the five case studies reported in the table is presented in detail in the five subsections below, followed by a sixth subsection on uncertainty of the multi-fidelity estimates. For all problems, we use a normalized graph Laplacian with $p=q=\frac{1}{2}$.
@@
$e_{k}^{(i)}=\frac{|\bar{u}^{(i)}_{k}-\hat{u}^{(i)}_{k}|}{\frac{1}{N}\sum_{j=1}^{N}|\hat{u}^{(j)}_{k}|}\times 100\%,\qquad i\in\{1,\,\dots\,,N\}\quad k\in\{1,\,\dots,\,D\}.$ (5.1)
@@
$e^{(i)}=\frac{\|\bar{\bm{u}}^{(i)}-\hat{\bm{u}}^{(i)}\|_{2}}{\frac{1}{N}\sum_{j=1}^{N}\|\hat{\bm{u}}^{(j)}\|_{2}}\times 100\%,\qquad i\in\{1,\,\dots\,,N\}.$ (5.2)
@@
Problem Case 1 Case 2 Case 3 Case 4 Case 5
Physical model Elasticity Elasticity Darcy Beam Navier Stokes
Data dimension $D$ 5 10,201 10,201 512 221
Data type QoIs 2D Field 2D Field 1D Field 1D Field
LF data $N$ 6,000 3,000 6,000 10,000 10,000
HF data $M$ 150 (2.5%) 100 (3.3%) 120 (2%) 50 (0.5%) 100 (1%)
LF error % 6.82 29.49 25.2 30.7 18.63
MF error % 1.67 6.41 3.96 4.43 3.87
Error reduction 75.5% 78.3% 84.3% 85.6% 79.2%
Table 3: Summary of the attributes of the numerical experiments (top) and the performance of the multi-fidelity approach (bottom).
@@ 5.1 Force and traction attributes of an elastic body with a stiff inclusion
We consider a soft square sheet in plane stress with an internal stiffer elliptic inclusion. The length of the edge of the square is $L=10\,\mathrm{cm}$, and its Young’s modulus is $E=1\mathrm{MPa}$, whereas the Young’s modulus of the inclusion is a parameter. Both background and the inclusion are incompressible. The bottom edge of the square is fixed, while a uniform downward displacement of $v_{0}=-5\mathrm{mm}$ is applied to the top edge. The vertical edges are traction-free in both horizontal and vertical directions, and the top edge is traction free only in the horizontal direction (Figure 2). The objective is to predict attributes of the vertical traction field on the upper edge as a function of the shape, stiffness, orientation and location of the inclusion. This problem is described in detail in [42] and is motivated by the need to identify stiff tumors within a soft background tissue, which is particularly relevant to detecting and diagnosing breast cancer tumors [61, 62].
@@
[…]
@@
We employ two finite element method-based solvers differing in mesh density to solve the problem. The low-fidelity model uses a coarse mesh with 200 triangular elements, and the high-fidelity model uses a fine mesh with 20,000 elements. The solution for high-fidelity model is verified to be mesh converged.
@@
We sample $N=6,000$ instances of the input parameters from a uniform distribution and run the low- and high-fidelity models. Thereafter, we normalize the low- and high-fidelity quantities of interest using (2.1), so that each low-fidelity component has zero-mean and unit standard deviation. We then use $M=150$ high-fidelity data points ($2.5\%$ of the total) for generating the multi-fidelity data, and the remainder for testing the performance of the method.
@@
We quantify the accuracy of the low- and multi-fidelity data via the error defined in (5.1), and compare the distribution of these two errors for each component in Figure 4. For every component, the distribution of the error for the multi-fidelity data is closer to zero, and is narrower when compared with the distribution of the error for the low-fidelity data. In Table 5 we report the mean error across all points for each component separately, while in Table 3 we report the average of these errors. We observe that the multi-fidelity update has reduced the error in the low-fidelity data by 70–84%.
Quantity of interest $f_{1}$ $f_{2}$ $f_{3}$ $f_{4}$ $\sigma_{yy}^{\mathrm{max}}$
Low-fidelity error 5.74 4.8 7.57 4.93 11.04
Multi-fidelity error 0.92 1.45 1.66 0.96 3.36
Error reduction 84 % 69.8.1% 78.1% 80.5% 69.6%
Table 5: Error in the low- and multi-fidelity data for each output component for the elasticity problem (Case 1).
Figure 3: Low-, multi-, and high-fidelity datasets for the elasticity problem (Case 1). The low-fidelity data points are shown together with points where the high-fidelity data are sampled (shown by blue dots).
Figure 4: Error distribution for the low- and multi-fidelity data for each output component for the elasticity problem (Case 1).
@@ 5.2 Displacement field of an elastic body with a stiff inclusion
The input parameters for this problem include the orientation of the elliptical inclusion $\theta$, its Young’s Modulus $E_{I}$, its major semi-axis $a$, and the uniform vertical traction applied on the top edge $t_{y}$. The minor semi-axis is set to $b=\frac{1}{\pi a}$, to maintain a constant area for the elliptical inclusion. In Table 6 we provide the minimum and maximum values for these parameters. The output quantity of interest is the vertical displacement field $u_{y}(x,y)$ in the entire domain, sampled at $201101^{2}=10,$ 201 points on a uniform grid. We note that for every value of applied traction, $t_{y}$, the vertical displacement within the elastic body can be decomposed into a part that varies linearly from zero at $y=0$ to a maximum value at $y=1$ and another part that is nonlinear in $y$. For the linear part of the displacement the value at $y=1$ edge is set to the average vertical displacement at this edge. We also note that both the low- and high-fidelity models are able to capture the linear part accurately. Therefore, the main goal of the multi-fidelity approach is to use the high-fidelity data to improve the accuracy of the non-linear part of the vertical displacement. For this reason, for every realization of the input parameters, we compute the linear part of the displacement from the low-fidelity model and subtract it from the low-fidelity data and high-fidelity data.
@@
We use two finite element-based models to compute the low- and high-fidelity data. The high-fidelity model employs a structured triangular mesh with $20,000$ elements, while the low-fidelity model uses a much coarser structured triangular mesh with only $32$ elements. We interpolate the low- and high-fidelity solutions on to the same uniform grid to ensure that the dimensionality of these data is consistent.
@@
We consider $N=3,000$ pairs of low- and high-fidelity data, obtained by sampling the input parameters from a uniform distribution. Each low-fidelity data is scaled using (2.2), and corresponding high-fidelity data are scaled by the same procedure. We compute the multi-fidelity model with $M=100$ high-fidelity training data points ($3.3\%$ of the total). Figure 5 shows examples of the resulting low- and multi-fidelity fields, along with their differences relative to the high-fidelity solutions. We observe that error in the multi-fidelity field is much smaller.
We quantify the performance of the low- and multi-fidelity models via the error defined in (5.2). Figure 6 is a histogram of the distribution of these errors for the low- and multi-fidelity data. From this plot we observe that the multi-fidelity data has significantly lower errors, and the spread in the error is also smaller. The mean error for the low-fidelity data is 29.49%, whereas for the multi-fidelity data is 6.41%, which amounts to a 78.3% reduction (see Table 3).
Figure 5: Typical low- and multi-fidelity vertical displacement fields and corresponding errors for the elasticity problem (Case 2).
Figure 6: Error distribution for the low- and multi-fidelity data for the elasticity problem (Case 2).
@@ 5.3 Darcy flow
$-\nabla\cdot(\eta(\bm{x})\nabla p(\bm{x}))$ $=g(\bm{x}),\qquad\bm{x}\in\Omega$
$\int_{\Omega}p(\bm{x})\mathrm{d}\bm{x}=0.$
@@
$g(\bm{x})=\begin{cases}+1,&\text{if }x^{2}+(y-1)^{2}<0.15^{2},\\ -1,&\text{if }(x-1)^{2}+y^{2}<0.15^{2},\\ 0,&\text{otherwise}.\end{cases}$
Figure 7: Schematic of the domain, source term and permeability field for the Darcy flow problem (Case 3).
@@
$\eta(\bm{x})$ $=\exp\left({G(\bm{x})}\right),\;G(\cdot)\sim\mathcal{N}(0,k(\cdot,\,\cdot))$
$k(\bm{x},\,\bm{x}^{\prime})$ $=\exp\left(-\frac{\|\bm{x}-\bm{x}^{\prime}\|^{2}_{2}}{(0.25)^{2}}\right).$
@@
We compute both low- and high-fidelity solutions using finite element-based models. The high-fidelity model employs a mesh of $20,000$ triangular elements, while the low-fidelity model uses a mesh with $72$ triangular elements. We interpolate both low- and high-fidelity solutions onto a uniform grid of $201101^{2}=10,$ 201 points to ensure they have the same dimension.
@@
The input to the problem is an instance of the permeability field $\eta^{(i)}(\bm{x})$. We generate this field efficiently by utilizing a truncated Karhunen–Loève expansion [63, 64], retaining the first 200 terms. The quantity of interest is the resulting pressure field $p^{(i)}(\bm{x})$ defined over the whole domain on a uniform grid of $101\times 101$ points. Therefore, for this problem $D=10,201$.
@@
The low-fidelity dataset consists of $N=6,000$ data points, and we use $M=120$ high-fidelity training points to compute the multi-fidelity estimates.
In Figure 8, we have plotted four instances of the input permeability, and the corresponding low-and multi-fidelity pressure fields normalized with Eq. (2.2), as well as the difference between these fields and their high-fidelity counterpart. We observe that while for all cases the pressure field varies from a large value on the top-left corner to a small value on the bottom-right corner, the permeability field has a significant effect on this variation. Generally speaking, large permeability values lead to a uniform distribution of pressure, while small values lead to sharper changes in pressure. The low- and multi-fidelity pressure fields are qualitatively similar, however on closer look there are discernible differences. This is made clear by comparing the error fields for the two, which are plotted on the same scale. The error in the low-fidelity field is much larger when compared with the multi-fidelity field.
In Figure 9 we have plotted the histogram for the error distributions (as defined in Eq. (5.2)) for the low-and multi-fidelity data. Once again we observe that the multi-fidelity data has much smaller error and that this error is more tightly centered about its mean. The mean values of the low- and multi-fidelity errors are 25.2% and 3.96%, respectively (see Table 3), which implies a 84.3% percentage reduction in error due to the multi-fidelity approach.
Figure 8: Instances of permeability fields, low- and multi-fidelity solutions and their errors for the Darcy flow problem (Case 3).
Figure 9: Histogram of the error distribution for the low- and multi-fidelity data for the Darcy flow problem (Case 3).

### v2

@@ 5 Numerical results
In this section we quantify the performance of the multi-fidelity approach for a suite of diverse problems in computational physics. These problems are governed by different physical models that include the equations of linear elasticity, Darcy’s flow, Euler–Bernoulli beam theory, and the incompressible Navier Stokes equations. The data types considered include vectors of quantities of interest (QoIs), and one- and two-dimensional fields, where the dimension of the data space, $D$, ranges from $5$ to $10,201$. The size of the low-fidelity datasets ranges from $3,000$ to $10,000$, and in each case the fraction of the number of high-fidelity data points to the number of low-fidelity data points is small; it lies between $0.5\%$ and $3.3\%$. These attributes are summarized in Table 3. Each of the five case studies reported in the table is presented in detail in the five subsections below, followed by a sixth subsection on uncertainty of the multi-fidelity estimates. For all problems, we use a normalized graph Laplacian with $p=q=\frac{1}{2}$.
For all of the experiments, $N$ is sufficiently small as to not require the numerical approximations described in Section 3, and therefore in all of the below we will take $K=N$, i.e., we will use all of the eigenvectors of the graph Laplacian. In forthcoming work we will explore the effect of varying $K$ on the multi-fidelity accuracy.
@@
$e_{k}^{(i)}=\frac{|\bar{u}^{(i)}_{k}-\hat{u}^{(i)}_{k}|}{\frac{1}{N}\sum_{j=1}^{N}|\hat{u}^{(j)}_{k}|}\times 100\%,\qquad i\in\{1,\,\dots\,,N\},\quad k\in\{1,\,\dots,\,D\}.$ (64)
@@
$e^{(i)}=\frac{\|\bar{\bm{u}}^{(i)}-\hat{\bm{u}}^{(i)}\|_{2}}{\frac{1}{N}\sum_{j=1}^{N}\|\hat{\bm{u}}^{(j)}\|_{2}}\times 100\%,\qquad i\in\{1,\,\dots\,,N\}.$ (65)
@@
Problem Case 1 Case 2 Case 3 Case 4 Case 5 Physical model Elasticity Elasticity Darcy Beam Navier Stokes Data dimension $D$ 55 10,20110{,}201 10,20110{,}201 $512$ $221$ Data type QoIs 2D Field 2D Field 1D Field 1D Field LF data NN 6,0006{,}000 3,0003{,}000 6,0006{,}000 10,00010{,}000 10,00010{,}000 HF data $M$ $150$ ($2.5\%$) $100$ ($3.3\%$) $120$ ($2\%$) $50$ ($0.5\%$) $100$ ($1\%$) Cost ratio $\approx 1,000$ $\approx 15,000$ $\approx 5,000$ N/A $\approx 9,400$ [CHENG2024116793] LF error $6.82$ ($5.11$) $29.49$ ($12.19$) $25.2$ ($5.4$) $30.7$ ($2.68$) $18.63$ ($5.96$) MF error 1.67 ($2.09$) 6.41 ($4.77$) 3.96 ($2.18$) 4.43 ($3.34$) 3.87 ($2.32$) Error reduction $75.5\%$ $78.3\%$ $84.3\%$ $85.6\%$ $79.2\%$
Table 3: Summary of the attributes of the numerical experiments (top) and the accuracy of the low- and multi-fidelity models (bottom). In particular, we list the data dimensions, the datasets size, the ratio of HF data used, and an estimate of the ratio between the computational cost of the high- and low-fidelity models. For the beam problem, the low-fidelity model is an analytical model with negligible cost. Finally, we report the mean and the standard deviation (in parenthesis) of the error distribution of the low- and multi-fidelity models.
@@ 5.1 Force and traction attributes of an elastic body with a stiff inclusion
We consider a soft square sheet in plane stress with an internal stiffer elliptic inclusion. The length of the edge of the square is $L=10\,\mathrm{cm}$, and its Young’s modulus is $E=1\mathrm{MPa}$, whereas the Young’s modulus of the inclusion is a parameter. Both background and the inclusion are incompressible. The bottom edge of the square is fixed, while a uniform downward displacement of $v_{0}=-5\mathrm{mm}$ is applied to the top edge. The vertical edges are traction-free in both horizontal and vertical directions, and the top edge is traction free only in the horizontal direction (Figure 2). The objective is to predict attributes of the vertical traction field on the upper edge as a function of the shape, stiffness, orientation and location of the inclusion. This problem is described in detail in [pinti2023graph] and is motivated by the need to identify stiff tumors within a soft background tissue, which is particularly relevant to detecting and diagnosing breast cancer tumors [sarvazyan2012mechanical, barbone2010review].
@@
$f_{i}=\int_{(i-1)\frac{L}{4}}^{i\frac{L}{4}}\sigma_{yy}(x,\,L)\mathrm{d}x,\;\;\;\;i\in\{1,\,\dots,\,4\}.$ (66)
We also include the maximum value of vertical traction on the top edge as an additional feature, leading to $5$ quantities of interest: $u_{i}=f_{i}$, for each $i\in\{1,\,\dots,\,4\}$, and $u_{5}=\max_{x}|\sigma_{yy}(x,\,L)|=:\sigma_{yy}^{\mathrm{max}}$. As the location, orientation and size of the inclusion is varied, the traction field on the top surface changes, which in turn changes the components of the localized forces, and the maximum value of traction.
@@
We employ two finite element method-based solvers differing in mesh density to solve the problem. The low-fidelity model uses a coarse mesh with 200 triangular elements, and the high-fidelity model uses a fine mesh with 20,000 elements. The solution for high-fidelity model is verified to be mesh converged. The ratio of the computational costs of the two models is a power of the mesh size ratio. The exact exponent lies between 1 and 2, depending on the solver and the problem. A typical value is approximately 1.5, leading to a computational cost ratio of around 1,000.
@@
We sample $N=$ 6,0006{,}000 instances of the input parameters from a uniform distribution and run the low- and high-fidelity models. Thereafter, we normalize the low- and high-fidelity quantities of interest using (1), so that each low-fidelity component has zero-mean and unit standard deviation. We then use $M=$ $150$ high-fidelity data points ($2.5\%$ of the total) for generating the multi-fidelity data, and the remainder for testing the performance of the method.
@@
We quantify the accuracy of the low- and multi-fidelity data via the error defined in (64), and compare the distribution of these two errors for each component in Figure 4. For every component, the distribution of the error for the multi-fidelity data is closer to zero, and is narrower when compared with the distribution of the error for the low-fidelity data. In Table 5 we report the mean and the standard deviation of the error distributions across all points for each component separately, while in Table 3 we report the average of these values. We observe that the multi-fidelity update has reduced the error in the low-fidelity data by 70–84%.
Quantity of interest $f_{1}$ $f_{2}$ $f_{3}$ $f_{4}$ $\sigma_{yy}^{\mathrm{max}}$ Low-fidelity error 5.74 (2.61) 4.8 (4.81) 7.57 (5.64) 4.93 (4.72) 11.04 (7.78) Multi-fidelity error 0.92 (0.97) 1.45 (1.86) 1.66 (1.85) 0.96 (1.14) 3.36 (4.62) Error reduction 84 % 69.8% 78.1% 80.5% 69.6%
Table 5: Error in the low- and multi-fidelity data for each output component for the elasticity problem (Case 1), together with its mean and standard deviation (in parenthesis).
Figure 3: The figure shows orthogonal projections of the datasets for the low-fidelity (left column), multi-fidelity (center column), and high-fidelity (right column) models for the elasticity problem (Case 1). Each row shows a different orthogonal plane. The low-fidelity data points, in column 1, exhibit greater range compared to the high-fidelity point cloud. We note how the multi-fidelity points show better agreement with the target high-fidelity distribution.
Figure 4: Histogram plots of the error distribution for the low- and multi-fidelity data for the elasticity problem (Case 1). Each plot shows the distribution of the error of the two models for one component of the vector of quantities of interest across the whole datasets. The values of mean and standard deviation of these distributions are reported in Table 5.
@@ 5.2 Displacement field of an elastic body with a stiff inclusion
The input parameters for this problem include the orientation of the elliptical inclusion $\theta$, its Young’s Modulus $E_{I}$, its major semi-axis $a$, and the uniform vertical traction applied on the top edge $t_{y}$. The minor semi-axis is set to $b=\frac{1}{\pi a}$, to maintain a constant area for the elliptical inclusion. In Table 6 we provide the minimum and maximum values for these parameters. The output quantity of interest is the vertical displacement field $u_{y}(x,y)$ in the entire domain, sampled at $101^{2}=10,201$ points on a uniform grid. We note that for every value of applied traction, $t_{y}$, the vertical displacement within the elastic body can be decomposed into a part that varies linearly from zero at $y=0$ to a maximum value at $y=1$ and another part that is nonlinear in $y$. For the linear part of the displacement the value at $y=1$ edge is set to the average vertical displacement at this edge. We also note that both the low- and high-fidelity models are able to capture the linear part accurately. Therefore, the main goal of the multi-fidelity approach is to use the high-fidelity data to improve the accuracy of the non-linear part of the vertical displacement. For this reason, for every realization of the input parameters, we compute the linear part of the displacement from the low-fidelity model and subtract it from the low-fidelity data and high-fidelity data.
@@
We use two finite element-based models to compute the low- and high-fidelity data. The high-fidelity model employs a structured triangular mesh with $20,000$ elements, while the low-fidelity model uses a much coarser structured triangular mesh with only $32$ elements. Given the mesh size ratio of $\frac{20,000}{32}=625$, and assuming an exponent for the cost power law of 1.5, the ratio of the computational cost of the two models is approximately 15,000. We interpolate the low- and high-fidelity solutions on to the same $101\times 101$ uniform grid to ensure the dimensionality $D=$ 10,20110{,}201.
@@
We consider $N=$ 3,0003{,}000 pairs of low- and high-fidelity data, obtained by sampling the input parameters from a uniform distribution. Each low-fidelity data is scaled using (2), and corresponding high-fidelity data are scaled by the same procedure. We compute the multi-fidelity model with $M=$ $100$ high-fidelity training data points ($3.3\%$ of the total). Figure 5 shows examples of the resulting low- and multi-fidelity fields, along with their differences relative to the high-fidelity solutions. We observe that error in the multi-fidelity field is much smaller.
[…]
Figure 5: Typical instances of the non-linear part of the vertical displacement fields from low- and multi-fidelity models for the elasticity problem (Case 2). The figure includes plots of point-wise error between each model and the high-fidelity field.
Figure 6: Histogram of error distribution for low- and multi-fidelity datasets for the elasticity problem (Case 2). The error distribution for the low-fidelity data has a mean of 29.4929.49% and a standard deviation of $12.19$. In contrast, the multi-fidelity data shows a reduced mean error of 6.416.41% with a standard deviation of $4.77$.
@@ 5.3 Darcy flow
$-\nabla\cdot(\eta(\bm{x})\nabla p(\bm{x}))$ $=g(\bm{x}),\qquad\bm{x}\in\Omega$ (67)
$\int_{\Omega}p(\bm{x})\mathrm{d}\bm{x}=0.$ (68)
@@
$g(\bm{x})=\begin{cases}+1,&\text{if }x^{2}+(y-1)^{2}<0.15^{2},\\ -1,&\text{if }(x-1)^{2}+y^{2}<0.15^{2},\\ 0,&\text{otherwise}.\end{cases}$ (69)
Figure 7: Schematic of the domain, source term and permeability field for the Darcy flow problem (Case 3). A source term located in the top left corner injects fluid into the domain, while a sink in the bottom right corner allows the fluid to exit. The input parameter is an instance of permeability field $\eta(\bm{x})$, and the quantity of interest is the resulting pressure field $p(\bm{x})$ of the fluid percolating through the medium.
@@
$\eta(\bm{x})$ $=\exp\left({G(\bm{x})}\right),\;G(\cdot)\sim\mathcal{N}(0,k(\cdot,\,\cdot))$ (70)
$k(\bm{x},\,\bm{x}^{\prime})$ $=\exp\left(-\frac{\|\bm{x}-\bm{x}^{\prime}\|^{2}_{2}}{(0.25)^{2}}\right).$ (71)
@@
We compute both low- and high-fidelity solutions using finite element-based models. The high-fidelity model employs a mesh of $20,000$ triangular elements, while the low-fidelity model uses a mesh with $72$ triangular elements. With a mesh size ratio of $\frac{20,000}{72}=278$, and assuming an exponent for the cost power law of 1.5, the ratio of the computational cost of the two models is approximately 5,000. We interpolate both low- and high-fidelity solutions onto a uniform grid of $101^{2}=10,201$ points to ensure they have the same dimension.
@@
The input to the problem is an instance of the permeability field $\eta^{(i)}(\bm{x})$. We generate this field efficiently by utilizing a truncated Karhunen–Loève expansion [Huang-KL, Stefanou2007AssessmentOS], retaining the first 200 terms. The quantity of interest is the resulting pressure field $p^{(i)}(\bm{x})$ defined over the whole domain on a uniform grid of $101\times 101$ points. Therefore, for this problem $D=$ 10,20110{,}201.
@@
The low-fidelity dataset consists of $N=$ 6,0006{,}000 data points, and we use $M=$ $120$ high-fidelity training points ($2\%$ of the total) to compute the multi-fidelity estimates.
In Figure 8, we have plotted four instances of the input permeability, and the corresponding low-and multi-fidelity pressure fields normalized with Eq. (2), as well as the difference between these fields and their high-fidelity counterpart. We observe that while for all cases the pressure field varies from a large value on the top-left corner to a small value on the bottom-right corner, the permeability field has a significant effect on this variation. Generally speaking, large permeability values lead to a uniform distribution of pressure, while small values lead to sharper changes in pressure. The low- and multi-fidelity pressure fields are qualitatively similar, however on closer look there are discernible differences. This is made clear by comparing the error fields for the two, which are plotted on the same scale. The error in the low-fidelity field is much larger when compared with the multi-fidelity field.
[…]
Figure 8: Instances of permeability fields and corresponding low- and multi-fidelity pressure distributions for the Darcy flow problem (Case 3). Plots of point-wise error fields with respect to the high-fidelity field are also shown.
Figure 9: Histogram of error distribution for low- and multi-fidelity datasets for the Darcy flow problem (Case 3). The mean and standard deviation of the low-fidelity distribution are 25.225.2% and $5.4$, respectively. On the other hand, the mean and standard deviation of the distribution for the multi-fidelity data are 3.963.96% and $2.18$, respectively.